# Heterogeneous Treatment Effects and

### v1

@@ 4.1 HTEs and Mechanisms with Directly-Affected Outcomes
Collectively, the concept of MIVs and the two exclusion assumptions convey the basic intuition that our initial results draw upon. For HTEs to provide information about a mechanism, it must be the case that the moderator of interest affects: (1) the degree to which treatment activates a mechanism or (2) the mechanism’s effect on our outcome of interest, $Y$. Recognition of the former scenario—consistent with our stylized example of voter updating—is well-known. The latter opens new possibilities for identifying moderators. The exclusion assumptions suggest that if a covariate, $X_{k}$, is a MIV for multiple mechanisms (or a direct effect), it cannot be used to confirm the presence of a given mechanism. This logic is straightforward, but it remains implicit in most uses of heterogeneous treatment effects to detect mechanisms. These intuitions give rise to Proposition 1.
@@
Suppose that $Y$ is directly affected by mechanism $j$ and Assumptions 1 and 2 hold with respect to $X_{k}$. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}_{j}^{MIV}$ for mechanism $j$.
Proposition 1 conveys important implications about our ability to use heterogeneous treatment effects to provide evidence for a mechanism. Recall that if $X_{k}$ is a MIV for mechanism $j$, $AIE_{j}(X_{k}=x)\neq AIE_{j}(X_{k}=x^{\prime})$ for some $x,x^{\prime}\in X_{k}$. This provides evidence that mechanism $j$ is active for at least one unit. The exclusion assumptions rule out the possibility that $AIE_{\neg j}(X_{k}=x)\neq AIE_{\neg j}(X_{k}=x^{\prime})$ for all other mechanisms ($\neg j$), as well as the possibility that $ADE(X_{k}=x)\neq ADE(X_{k}=x^{\prime})$. When these assumptions hold, HTEs in $X_{k}$ are thus sufficient to show that $X_{k}$ is a MIV for mechanism $j$. This conforms to standard interpretations that the presence of HTEs support arguments about mechanism activation. Nevertheless, this interpretation invokes the two exclusion assumptions, which are generally not invoked explicitly. We now consider the converse: the case when there exist no HTEs with respect to $X_{k}$.
@@
    $X_{k}\notin\textbf{X}_{j}^{MIV}$ for mechanism $j$.
@@
    No MIV exists.
Proposition 2 shows that a lack of HTEs provides less information with regard to mechanism activation than is generally asserted. Under the exclusion assumptions, there are two reasons why HTEs may not exist with respect to a covariate, $X_{k}$. First, it may be the case that $X_{k}$ is not a MIV for mechanism $j$. In this sense, we have misspecified the theoretical relationship between a given covariate and a mechanism. Second, it may be the case that no MIV exists for mechanism $j$. As we discuss in Corollary 1, there are two possible reasons why a MIV would not exist for mechanism $j$. Importantly, we show that this could happen with an active or an inert mechanism $j$.
@@
If no MIV exists for a mechanism $j$, there are two possibilities:
(1) Mechanism $j$ is not active.
(2) Mechanism $j$ is active, but there exists no $X$ for which $AIE_{j}(X=x)\neq AIE_{j}(X=x^{\prime})$.
Case (1) of Corollary 1 is implied by the definition of MIV. If a mechanism is inert—thereby producing an indirect effect of zero for all units—there cannot exist any MIVs, measured or unmeasured. In contrast, in Case (2), a mechanism can be active and produce the same indirect effect for all units. In this case, there are no covariates that moderate the indirect effect. These results show that, in contrast to standard interpretation, a lack of heterogeneity cannot tell us about whether a mechanism is active. Moreover, our theory could be misspecified, meaning that our postulated MIV, $X_{k}$ is not actually a MIV. An assessment of HTEs with respect to a single moderator cannot distinguish between these three possibilities. Nor can we assign probabilities to these (non-mutually exclusive) explanations.
Comparing Propositions 1 and 2, under the two exclusion assumptions, the presence of HTEs provides more information with regard to mechanisms than does the absence of HTE. In this sense, relying upon the a lack of heterogeneity to “rule out” a potential mechanism requires much stronger theoretical assumptions than is generally acknowledged. Specifically, we would need to assume that $X_{k}\in\textbf{X}_{j}^{MIV}$ in order to rule out the activation of mechanism $j$ [^17] Note that if we assume that $X_{k}\in\textbf{X}_{j}^{MIV}$ we have also implicitly assumed that there exists an $X_{j}$ for which $AIE_{j}(X_{j}=x)\neq AIE_{j}(X_{j}=x^{\prime})$. Indeed, such an assumption is precisely what we are trying to learn from the presence of HTEs in Proposition 1.
@@ 5 Indirectly-Affected Outcomes
When outcomes are directly affected by one or more mechanisms under a theoretical model, the existance of HTEs provides evidence that a mechanism is active, as we have shown in Proposition 1. Yet, the situation is more complicated in the case of indirectly-affected outcomes. As in our discussion of the exogenous shocks and voting, we observed heterogeneous treatment effects in both prior beliefs (the mechanism in the model) and in valence (not the mechanism in the model). In this section, we show that this finding is general to situations in which indirectly-affected outcomes are generated by a non-linear transformation of the relevant directly-affected outcome, as noted in Definition 3.
@@
$h(Y)$ =$=\begin{cases}1&Y\in(-\infty,c_{1}]\\ 2&Y\in(c_{1},c_{2}]\\ \vdots&\\ Q&Y\in(c_{Q+1},\infty),\end{cases}$ (15)
in which $c_{t}$ are increasing thresholds in the latent attitude. Our focus for this section is outcomes that are indirectly affected such that they are produced by a non-linear mapping from a directly affected outcome.
To understand what HTEs reveal with respect to an indirectly-affected outcome, it is useful to introduce one final concept. We will denote $\textbf{X}^{R}$ as the set of all possible pre-treatment covariates with non-zero effects on the directly-affected outcome, $Y$ [^18] Formally, if $X_{k}\in\textbf{X}^{R}$, then there exist $x\neq x^{\prime}\in X_{k}$ such that $Y(Z,M(Z;X_{k}=x,X_{\neg k});X_{k}=x,X_{\neg k})\neq Y(Z,M(Z;X_{k}=x^{\prime},X_{\neg k});X_{k}=x^{\prime},X_{\neg k})$. Covariates in $\textbf{X}^{R}$ can be thought of as “relevant” for predicting outcome $Y$. It is also useful to let X be the set of all possible pre-treatment covariates. It is clear that for any outcome, $Y$, and mechanism, $j$, $_{j}^{MIV}\subseteq\textbf{X}^{R}\subseteq\textbf{X}$ X. Typically, these subsets will be proper.
We now return to our main question of interest: what do HTEs reveal with regard to mechanisms? Proposition 3 considers the case when there are HTEs in a covariate $X_{k}$. Here, we can learn that $X_{k}\in\textbf{X}^{R}$, but this is not informative about whether $X\in\textbf{X}_{j}^{MIV}$, since $\textbf{X}_{j}^{MIV}\subseteq\textbf{X}^{R}$. In order to make an inference about mechanism $j$ (under the exclusion assumptions) we need to know whether $X_{k}\in\textbf{X}_{j}^{MIV}$. If $X_{k}\in\textbf{X}_{j}^{MIV}$ for mechanism $j$, then mechanism $j$ is active. If $X_{k}\notin\textbf{X}_{j}^{MIV}$ mechanism $j$ may or may not be active. The intuition for this result is straightforward. The non-linear mapping from $Y$ to $h(Y)$ “breaks” the additive separability between any $X\in\textbf{X}^{R}-\textbf{X}_{j}^{MIV}$ and the indirect effect of mechanism $j$, $IE_{j}$.
@@
It is useful to consider a simple numerical example of Proposition 3 where $X_{k}\in\textbf{X}^{R}-\textbf{X}_{j}^{MIV}$, meaning that $X_{k}\in\textbf{X}^{R}$ but $X_{k}\notin\textbf{X}_{j}^{MIV}$. Suppose that we are interested in how a mobilization treatment, $Z_{i}\in\{0,1\}$, affects citizens’ decision to vote. Consider two covariates that predict turnout: $X_{1}\sim\mathcal{N}(0,1)$ and $X_{2}\in\{0,1\}$. We will further assume that $X_{1}$ is a MIV for the unique mechanism, such that:
$M_{1}(Z;\textbf{X})$ $=(1+Z)X_{1}$
@@
In this case, there is only one mechanism so Assumptions 1-2 hold by construction. Now, suppose that a researcher mistakenly thought that $X_{2}$ was a MIV for a mechanism (either $M_{1}$ or a non-existent mechanism). Since we have constructed the data generating process, we know that it is not: $X_{2}\in\textbf{X}^{R}-\textbf{X}_{j}^{MIV}$. Evaluating the CATE of $Z$ on turnout when $X_{2}=1$, we have:
@@
Note that $\Phi(\cdot)$ is the cdf of the standard normal distribution, which we invoke because $X_{1}\sim\mathcal{N}(0,1)$. Using the same approach it is straightforward to see that $CATE(X_{2}=0)=\Phi(0)-\Phi(0)=0$. So it is clear that we have HTEs in $X_{2}$ because $CATE(X_{2}=1)\neq CATE(X_{2}=0)$. Remember that, by construction, $X_{2}$ is not a MIV, though it is relevant. This numerical example is analogous to the issue that arises with valence in the analysis of our motivating example.
While it is straightforward to construct such many such examples mathematically, it is reasonable to ask: is this plausible in “real world” examples? We cannot answer this question empirically without observing which mechanisms are at work. Indeed, this is the whole problem that quantitative analysis of mechanisms seeks to (indirectly) answer! However, we can conduct simulations that use real data. In G, we use 2020 ANES data in a Monte Carlo simulation motivated by the theoretical model of Little, Schnakenberg, and Turner (2022). This paper proposes two mechanisms that account for how citizens update beliefs in response to new information: accuracy and directional motives. Their model implies that partisanship (or ideology) should be a MIV for directional but not accuracy motives, which would allow researchers to look for heterogeneity in partisanship to assess the presence of directional motives as a mechanism.
Following this logic, we simulate different treatment effects on a measure of latent attitudes about greenhouse gas regulation. Some simulations allow for directional motives, others shut down this channel. We show that when treatment has a non-zero effect on the latent attitudes for any subset of partisans, as we would expect from accuracy motives alone, there are HTEs in partisanship on a binary measure of preferences for stronger greenhouse gas regulation. This means that we cannot use these HTEs to assess the presence of directional motives with this indirectly measured outcome. Further, using the ANES sample, for some simulated effect sizes, there is greater statistical power to detect heterogeneity in partisanship when effects on the latent variable are homogeneous than when they are heterogeneous.
We now return to a final case of our theoretical analysis by asking when we are examining an outcome that may be indirectly affected by mechanism $j$, what can we learn from a lack of HTEs? Proposition 4 indicates that in this case, we can infer that $_{k}\in\textbf{X}$ X. This is obviously a vacuous result. We already know that $_{k}\in\textbf{X}$ X since $X_{k}$ is a covariate and X is the set of all covariates. We purposely state a vacuous result to emphasize how little can be ascertained about mechanisms from the non-existence of HTEs when outcomes are indirectly affected by a mechanism.
@@
Often we make assumptions about the mapping $h$. For example, the mapping in (15) imposes assumptions about how latent attitudes translate into Likert-scale responses. When we are willing to make such assumptions, we can refine Proposition 4 slightly. Specifically, in Proposition 5, we show that under Assumptions 1-2, for absolutely continuous directly-affected outcome, $Y$ and the non-linear transformation in (15) (for any $Q\geq 2$ categories), if HTEs do not exist with respect to $X_{k}$, then $X_{k}\notin\textbf{X}^{R}$. Because $\textbf{X}_{j}^{MIV}\subseteq\textbf{X}^{R}$ we know then that $X_{k}\notin\textbf{X}_{j}^{MIV}$ if $X_{k}\notin\textbf{X}^{R}$. But as in Proposition 2 and Corollary 1, there are multiple possible explanations: our theory about how $X_{k}$ relates to mechanism $j$ could be wrong or no MIV exists for mechanism $j$. These possibilities mean that we cannot make an inference about mechanism (non)-activation from the absence of HTEs with respect to $X_{k}$.
@@
Directly affected by mechanism $j$ Indirectly affected by mechanism $j$
$\exists$ HTEs wrt $X_{k}$: $X_{k}\in\textbf{X}_{j}^{MIV}$ $X_{k}\in\textbf{X}^{R}$
$\implies M_{j}$ is active. $M_{j}$ is active or inactive
$\nexists$ HTEs wrt $X_{k}$: $X_{k}\notin\textbf{X}_{j}^{MIV}$ and/or $\textbf{X}_{j}^{MIV}=\emptyset$ $_{k}\in\textbf{X}$ X (vacuous)
$M_{j}$ is active or inactive $M_{j}$ is active or inactive
Table 2: In this table, $\textbf{X}_{j}^{MIV}$ is defined with respect to mechanism $j$. The results in each cell invoke Assumptions 1 and 2.

### v2

@@ 4.1 HTEs and Mechanisms with Directly-Affected Outcomes
Collectively, the concept of MDVs and the exclusion assumption convey the basic intuition that our initial results draw upon. For HTEs to provide information about a mechanism, it must be the case that the moderator of interest affects: (1) the degree to which treatment activates a mechanism or (2) the mechanism’s effect on our outcome of interest, $Y$. Recognition of the former scenario—consistent with our stylized example of voter updating—is well-known. The latter opens new possibilities for identifying moderators. The exclusion assumption suggests that if a covariate, $X_{k}$, is a MDV for direct effects (or multiple mechanisms), it cannot be used to confirm the presence of a given mechanism. This logic is straightforward, but it remains implicit in most uses of heterogeneous treatment effects to detect mechanisms. These intuitions give rise to Proposition 1.
@@
Suppose that $Y$ is directly affected by mechanism $M$ and Assumption 2 holds with respect to $X_{k}$. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}^{MDV}$ for mechanism $M$.
Proposition 1 conveys important implications about our ability to use heterogeneous treatment effects to provide evidence for a mechanism. Recall that if $X_{k}$ is a MDV for mechanism $M$, $AIE(X_{k}=x)\neq AIE(X_{k}=x^{\prime})$ for some $x,x^{\prime}\in X_{k}$. This provides evidence that mechanism $M$ is active for at least one unit. The exclusion assumption rules out the possibility that $ADE(X_{k}=x)\neq ADE(X_{k}=x^{\prime})$. When the assumption holds, HTEs in $X_{k}$ are thus sufficient to show that $X_{k}$ is a MDV for mechanism $M$. This conforms to standard interpretations that the presence of HTEs support arguments about mechanism activation. Nevertheless, this interpretation invokes the exclusion assumption, which is generally not invoked explicitly. We now consider the converse: the case when there exist no HTEs with respect to $X_{k}$.
@@
Suppose that $Y$ is directly affected by mechanism $M$ and Assumptions 1 and 2 hold. If no HTEs exist with respect to $X_{k}$, at least one of the following must be true:
@@
    $X_{k}\notin\textbf{X}^{MDV}$ for mechanism $M$.
@@
    No MDV exists.
Proposition 2 shows that a lack of HTEs provides less information with regard to mechanism activation than is generally asserted. Under the exclusion assumption, there are two reasons why HTEs may not exist with respect to a covariate, $X_{k}$. First, it may be the case that $X_{k}$ is not a MDV for mechanism $M$. In this sense, we have misspecified the theoretical relationship between a given covariate and a mechanism. Second, it may be the case that no MDV exists for mechanism $M$. As we discuss in Corollary 1, there are two possible reasons why a MDV would not exist for mechanism $M$. Importantly, we show that this could happen with an active or an inert mechanism $M$.
@@
If no MDV exists for a mechanism $M$, there are two possibilities:
(1) Mechanism $M$ is not active.
(2) Mechanism $M$ is active, but there exists no $X$ for which $AIE(X=x)\neq AIE(X=x^{\prime})$.
Case (1) of Corollary 1 is implied by the definition of MDV. If a mechanism is inert—thereby producing an indirect effect of zero for all units—there cannot exist any MDVs, measured or unmeasured. In contrast, in Case (2), a mechanism can be active and produce the same indirect effect for all units. In this case, there are no covariates that moderate the indirect effect. These results show that, in contrast to standard interpretation, a lack of heterogeneity cannot tell us about whether a mechanism is active. Moreover, our theory could be misspecified, meaning that our postulated MDV, $X_{k}$ is not actually a MDV. An assessment of HTEs with respect to a single moderator cannot distinguish between these three possibilities. Nor can we assign probabilities to these (non-mutually exclusive) explanations.
Comparing Propositions 1 and 2, under the exclusion assumption, the presence of HTEs provides more information with regard to mechanisms than does the absence of HTE. In this sense, relying upon the a lack of heterogeneity to “rule out” a potential mechanism requires much stronger theoretical assumptions than is generally acknowledged. Specifically, we would need to assume that $X_{k}\in\textbf{X}^{MDV}$ in order to rule out the activation of mechanism $M$. Indeed, such an assumption is precisely what we are trying to learn from the presence of HTEs in Proposition 1.
@@ 5 Indirectly-Affected Outcomes
When outcomes are directly affected by a mechanism under a theoretical model, the existence of HTEs provides evidence that a mechanism is active, as we have shown in Proposition 1. Yet, the situation is more complicated in the case of indirectly-affected outcomes. As in our discussion of the exogenous shocks and voting, for the vote choice outcome, we observed heterogeneous treatment effects in both prior beliefs (the mechanism in the model) and in valence (not the mechanism in the model). In this section, we show that this finding is general to situations in which indirectly-affected outcomes are generated by a non-linear transformation of the relevant directly-affected outcome, as noted in Definition 3.
@@
$h(Y)$ =$=\begin{cases}1&Y\in(-\infty,c_{1}]\\ 2&Y\in(c_{1},c_{2}]\\ \vdots&\\ Q&Y\in(c_{Q-1},\infty),\end{cases}$ (15)
in which $c_{t}$ are increasing thresholds in a latent attitude. Our focus for this section is outcomes that are indirectly affected such that they are produced by a non-linear mapping from a directly affected outcome.
To understand what HTEs reveal with respect to an indirectly-affected outcome, it is useful to introduce one final concept. We will denote $\textbf{X}^{R}$ as the set of all possible pre-treatment covariates with non-zero effects on the directly-affected outcome, $Y$ [^14] Formally, if $X_{k}\in\textbf{X}^{R}$, then there exist $x\neq x^{\prime}\in X_{k}$ such that $Y(Z,M(Z;X_{k}=x,X_{\neg k});X_{k}=x,X_{\neg k})\neq Y(Z,M(Z;X_{k}=x^{\prime},X_{\neg k});X_{k}=x^{\prime},X_{\neg k})$. Covariates in $\textbf{X}^{R}$ can be thought of as “relevant” for predicting outcome $Y$. It is also useful to let X be the set of all possible pre-treatment covariates. It is clear that for any outcome, $Y$, and mechanism $M$, $^{MDV}\subseteq\textbf{X}^{R}\subseteq\textbf{X}$ X. Typically, these subsets will be proper.
We now return to our main question of interest: what do HTEs reveal with regard to mechanisms? Proposition 3 considers the case when there are HTEs in a covariate $X_{k}$. Here, we can learn that $X_{k}\in\textbf{X}^{R}$, but this is not informative about whether $X\in\textbf{X}^{MDV}$, since $\textbf{X}^{MDV}\subseteq\textbf{X}^{R}$. In order to make an inference about mechanism $M$ (under the exclusion assumption) we need to know whether $X_{k}\in\textbf{X}^{MDV}$. If $X_{k}\in\textbf{X}^{MDV}$ for mechanism $M$, then mechanism $M$ is active. If $X_{k}\notin\textbf{X}^{MDV}$, mechanism $M$ may or may not be active. The intuition for this result is straightforward. The non-linear mapping from $Y$ to $h(Y)$ “breaks” the additive separability between any $X_{k}\in\textbf{X}^{R}-\textbf{X}^{MDV}$ and the indirect effect.
@@
Suppose that observed outcome $h(Y)$ is a non-linear mapping of directly-affected outcome $Y$ and Assumptions 1holds. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}^{R}.$
It is useful to consider a simple numerical example of Proposition 3 where $X_{k}\in\textbf{X}^{R}-\textbf{X}^{MDV}$, meaning that $X_{k}\in\textbf{X}^{R}$ but $X_{k}\notin\textbf{X}^{MDV}$. Suppose that we are interested in how a mobilization treatment, $Z_{i}\in\{0,1\}$, affects citizens’ decision to vote. Consider two covariates that predict turnout: $X_{1}\sim\mathcal{N}(0,1)$ and $X_{2}\in\{0,1\}$. We will further assume that $X_{1}$ is a MDV for the unique mechanism, such that:
$M(Z;\textbf{X})$ $=(1+Z)X_{1}$
@@
In this case, assumption 2 holds by construction. Now, suppose that a researcher mistakenly thought that $X_{2}$ was a MDV for a mechanism (either $M$ or a non-existent mechanism). Since we have constructed the data generating process, we know that it is not: $X_{2}\in\textbf{X}^{R}-\textbf{X}^{MDV}$. Evaluating the CATE of $Z$ on turnout when $X_{2}=1$, we have:
@@
Note that $\Phi(\cdot)$ is the cdf of the standard normal distribution, which we invoke because $X_{1}\sim\mathcal{N}(0,1)$. Using the same approach it is straightforward to see that $CATE(X_{2}=0)=\Phi(0)-\Phi(0)=0$. So it is clear that we have HTEs in $X_{2}$ because $CATE(X_{2}=1)\neq CATE(X_{2}=0)$. Remember that, by construction, $X_{2}$ is not a MDV, though it is relevant. This numerical example is analogous to the issue that arises with valence in the analysis of our motivating example.
While it is straightforward to construct infinte such examples numerically, it is reasonable to ask: is this plausible in “real world” examples? We cannot answer this question empirically without observing which mechanisms are at work. Indeed, this is the whole problem that quantitative analysis of mechanisms seeks to (indirectly) answer! However, we can conduct simulations that use real data. In H, we use 2020 ANES data in a Monte Carlo simulation motivated by the theoretical model of Little, Schnakenberg, and Turner (2022). This paper proposes two mechanisms that account for how citizens update beliefs in response to new information: accuracy and directional motives. Their model implies that partisanship (or ideology) should be a MDV for directional but not accuracy motives, which would allow researchers to look for heterogeneity in partisanship to assess the presence of directional motives as a mechanism.
Following this logic, we simulate different treatment effects on a measure of latent attitudes about greenhouse gas regulation. Some simulations allow for directional motives, others shut down this channel. We show that when treatment has a non-zero effect on the latent attitudes for any subset of partisans, as we would expect from accuracy motives alone, there are HTEs in partisanship on a binary measure of preferences for stronger greenhouse gas regulation. This means that we cannot use these HTEs to assess the presence of directional motives with this indirectly measured outcome. Further, using the ANES sample, for some simulated effect sizes, there is greater statistical power to detect heterogeneity in partisanship when simulated effects on the latent variable are homogeneou in partisanship than when they are heterogeneous.
We now return to a final case of our theoretical analysis by asking when we are examining an outcome that may be indirectly affected by mechanism $M$, what can we learn from a lack of HTEs? Proposition 4 indicates that in this case, we can infer that $_{k}\in\textbf{X}$ X. This is obviously a vacuous result. We already know that $_{k}\in\textbf{X}$ X since $X_{k}$ is a covariate and X is the set of all covariates. We purposely state a vacuous result to emphasize how little can be ascertained about mechanisms from the non-existence of HTEs when outcomes are indirectly affected by a mechanism.
@@
Suppose that observed outcome $h(Y)$ is a non-linear mapping of directly-affected outcome $Y$ and Assumptions 1 holds. If HTEs do not exist with respect to $X_{k}$, then $_{k}\in\textbf{X}.$ X .
@@
Often we make assumptions about the mapping $h$. For example, the mapping in (15) imposes assumptions about how latent attitudes translate into Likert-scale responses. When we are willing to make such assumptions, we can refine Proposition 4 slightly. Specifically, in Proposition 9, we show that under the exclusion assumption, for absolutely continuous directly-affected outcome, $Y$ and the non-linear transformation in (15) (for any $Q\geq 2$ categories), if HTEs do not exist with respect to $X_{k}$, then $X_{k}\notin\textbf{X}^{R}$. Because $\textbf{X}^{MDV}\subseteq\textbf{X}^{R}$ we know then that $X_{k}\notin\textbf{X}^{MDV}$ if $X_{k}\notin\textbf{X}^{R}$. But as in Proposition 2 and Corollary 1, there are multiple possible explanations: our theory about how $X_{k}$ relates to mechanism $M$ could be wrong or no MDV exists for mechanism $M$. These possibilities mean that we cannot make an inference about mechanism (non)-activation from the absence of HTEs with respect to $X_{k}$.
@@
Directly affected by mechanism $M$ Indirectly affected by mechanism $M$
$\exists$ HTEs wrt $X_{k}$: $X_{k}\in\textbf{X}^{MDV}$ $X_{k}\in\textbf{X}^{R}$
$\implies M$ is active. $M$ is active or inactive
$\nexists$ HTEs wrt $X_{k}$: $X_{k}\notin\textbf{X}^{MDV}$ and/or $\textbf{X}^{MDV}=\emptyset$ $_{k}\in\textbf{X}$ X (vacuous)
$M$ is active or inactive $M$ is active or inactive
Table 2: In this table, $\textbf{X}^{MDV}$ is defined with respect to mechanism $M$. The results in each cell invoke Assumption 2.